# Classification Diffusion Models: Revitalizing Density Ratio Estimation

### v2

@@ 4 Experiments
We train several models using CDM on two common datasets. For CIFAR-10 [26] we train both a class conditional model and an unconditional model. We also train a similar model for CelebA [30], using face images of size $64\times 64$. In Sec. 4.1, we demonstrate why DRE methods have failed in settings of complex high-dimensional data like images, and show how the incorporation of the MSE loss in our method, according to Theorem 3.1, overcomes these challenges. In Sec. 4.2, we compare our method with pre-trained DDMs with similar architectures, to disentangle the benefits of our method from other variables. We evaluate the performance of CDM as a denoiser, assess its generation quality using FID [17], and measure its likelihood modeling capabilities using NLL. Finally, in Sec. 4.3, we demonstrate the use of a different scheduler specifically tuned for DRE tasks. We show that incorporating this scheduler into our method achieves state-of-the-art NLL results among methods capable of outputting the likelihood in a single forward pass.
@@ 4.1 The Importance of Using Both Losses for Achieving an Optimal Classifier
In Theorem 3.1, we established that the MMSE denoiser corresponds to the gradient of the optimal classifier. Therefore, in theory it should be sufficient to optimize only one of the objectives, i.e. either use only a CE loss on the model’s output or use only an MSE loss on its derivative. In Table 1 we report the MSE, CE and classification accuracy achieved by models trained with different losses. We emphasize that the model trained using only MSE in this comparison, is a CDM model trained using Algorithm 1, and is not equivalent to a DDM model trained following (5). As evident from Table 1, when using only the CE loss, the MSE is high, and when using only the MSE loss the CE and classification accuracy are poor. An important point to notice is that even when training with the CE loss, the classifier’s accuracy is low. This is a key prerequisite for making DRE methods work. Specifically, as shown in [34], to avoid the density-chasm problem, the classification problem should be sufficiently hard, otherwise the classifier can easily discriminate between the classes even without having learned the correct density ratio. Next, we demonstrate that simply making the classification problem harder is still insufficient for learning the correct probability $p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})$ for any $t$.
Figure 3 shows the logits ($f_{\theta}({\bm{x}}_{t})$) for noisy images with different noise levels, comparing a model trained using CE to a model trained with both CE and MSE. As the results in Fig. 3 reflect, in both scenarios the prediction near the true label is the same, namely the CE works well in the local area of the correct noise level. However, the model trained without MSE displays significantly higher predicted logits for more distant noise levels compared to its counterpart model. Moreover, the logits of the model trained without MSE do not decrease monotonically as the distance from the actual noise level increases, in contrast to the expected behavior. This demonstrates the importance of the MSE loss for obtaining good prediction globally. As can be seen in Theorem 3.1, the denoiser at timestep $t$ depends on the predictions of the classifier in both $t$-th and $(T+1)$-th entries. Therefore, the addition of the MSE loss enforces the classifier to achieve accurate predictions in both entries, thereby ensuring accurate predictions globally.
@@ 4.2 Denoising Results, Image Quality and Negative Log Likelihood
We compare our method with pre-trained DDMs with similar architectures. Since CDM is a classifier and DDM is a denoiser, for CDM we alter the last two layers such that the output is a vector of logits. We additionally remove all its timestep conditioning layers. These changes have a negligible impact on the number of parameters in the model. For more details please refer to App. C.
@@
For image denoising, CDM surpasses pre-trained DDMs at high noise levels in terms of MSE, while achieving comparable MSE at lower noise levels as shown in Fig. 5. These quantitative results are corroborated by the qualitative examples in Fig. 5, which showcase image denoising results across various noise levels.
@@
Beyond denoising, CDM demonstrates the ability to generate high-quality, diverse images from CelebA and unconditional CIFAR-10 as illustrated in Fig. 1. We compare the quality of images generated using CDM to images sampled from pre-trained DDMs trained according to [19]. Image quality is assessed using 50k FID [17] with the train-set. For both CDMs and DDMs, we compare images sampled using several schedulers; DDPM [19], DDIM [38], and DPM-Solver (DPMS) [31] samplers, set with 1000, 50, and 25 timesteps for sampling, respectively. The results, shown in Table 2, demonstrate that CDM is at least comparable to pre-trained DDMs in image quality, outperforming them in most cases.
@@
Figure 4: MSE-based denoising results across different noise levels (timesteps $t$). In terms of MSE, our method significantly outperforms a pre-trained DDM with the same architecture, at high noise levels, while demonstrating comparable performance at lower noise levels.
@@
Figure 5: Denoising results for CelebA $\mathbf{64\times 64}$ at various noise levels. We show a comparison between a CDM and a pre-trained DDM for different noise levels. The right column shows CDM’s and DDM’s predictions for pure Gaussian noise. When an MMSE denoiser is given a pure noise as an input, it should theoretically predict the expectation of the prior distribution. As observed, DDM outputs a highly noisy image, whereas CDM generates an image much closer to the mean of the dataset.
@@
Additionally, as Classifier-free guidance (CFG) has become a cornerstone of conditional generation with diffusion models [5], we evaluated CDM’s effectiveness in applying CFG during conditional sampling tasks. As expected, incorporating CFG improved image quality beyond unconditional generation, as reflected in the conditional CIFAR-10 FID results of Table 2. More details and qualitative results are provided in Appendix C.5. These results showcase the effectiveness of CDM for image generation, showing it to be equal or better than a similar DDM.
@@
Finally, we calculate log-likelihoods and compare our NLL results to recent leading methods in Table 3. CDM demonstrates comparable performance on NLL estimation for CIFAR-10 compared to DDMs. Notably, CDM stands out as a more efficient method than existing ones, requiring only a single forward pass for NLL computation. Table 3 also includes CDM(unif.), as outlined below in Sec. 4.3. This version of our model, which is more suited for DRE tasks, improves our NLL predictions and achieves state-of-the-art results among methods employing a single step.
4.3 Uniform Noise Scheduling for Better Likelihood Estimation
@@
To see the effect of using a timestep scheduler tuned for DRE tasks, we repeat the CIFAR-10 unconditional experiment, with a different noise scheduler. Following Rhodes et al. [34] we use the scheduler defined as $\sqrt{1-\bar{\alpha}_{t}}=\frac{t}{T+1}$ and choose $T=1000$ similarly to our previous experiments. Utilizing this scheduler, we achieve a state-of-the-art single-step NLL of 2.98 at the expense of a higher FID of 10.28, when using the DDIM sampler with 50 steps. This trade-off highlights that the scheduler optimal for learning the data distribution may not be ideal for sampling. Future research endeavors could explore schedulers aimed at further enhancing the NLL. Further analysis of the difference between the schedulers from a classification perspective can be found in E.2
@@ 5 Related Work
Using the concept of DRE for learning distributions of data was initially studied by Gutmann and Hyvärinen [15]. Their noise contrastive estimation (NCE) method approximates the ratio between the density of the data distribution and that of white Gaussian noise. However, it struggles in practical scenarios where the gap between these distributions is large, as is the case for natural images [34]. Conditional noise contrastive estimation (CNCE) [4] is a slightly improved version of NCE, in which the classification problem is designed to be harder. Specifically, CNCE is based on training a classifier to predict the order of a pair of samples with closer densities, e.g. achieved by pairing a data sample with its noisy version.
Telescoping density-ratio estimation (TRE), proposed by Rhodes et al. [34], avoids direct classification between data and noise. Instead, it uses a gradual transition between those two distributions, and trains a classifiers to distinguish between samples from every pair of adjacent densities. Such a classifier learns the ratio between adjacent distributions, and the overall ratio between the data and noise distributions is computed by multiplying all intermediate ratios.
Extending this concept, Choi et al. [7] took the notion further by expanding from a finite set of intermediate densities to an infinite continuum. This was accomplished by establishing a link between the estimation of density ratios for infinitesimally close distributions and the principles of score matching [23, 39, 41], and training a model to predict the time score $\frac{\partial}{\partial t}\log p_{{\mathbf{x}}_{t}}$. In contrast, we draw a different connection which shows that an MMSE denoiser can be obtained as the gradient of an optimal noise level classifier. Also, to obtain the log ratio between the target and reference distributions, Choi et al. [7] need to solve an integral over the time-score using an ODE solver, while in our method this ratio can be calculated in a single NFE.
@@
Table 2: Image generation quality. We compare the FID (lower is better) achieved by DDM and CDM using three sampling schemes for CelebA and CIFAR-10. For conditional CIFAR-10 we train a DDM ourselves, as no model in the original implementation [19] supports CFG.
@@
Table 3: NLL (bits/dim) calculated on CIFAR-10 test-set. For each model we specify the number of NFEs required for calculating the NLL. CDM achieves state-of-the-art NLL among methods that use a single NFE.
@@ 6 Discussion and Conclusion
We developed an analytical connection between an MSE-optimal denoiser for white Gaussian noise and a cross-entropy-optimal classifier for predicting the noise level. We used this connection to propose CDM – a DRE based generative technique that is based on a noise-level classifier. Importantly, our classifier is trained using both a classification loss (CE) and a regression loss (MSE). We showed that this key component is what sets CDM apart from existing DRE based methods, and makes it the first instance of a DRE-based technique that can successfully generate images beyond MNIST.
Our approach is not free of limitations. A key challenge is that CDMs can be more computationally expensive than DDMs. Indeed, while DDMs require a single forward pass for each denoising step, CDMs requires both a forward pass and a backward pass. Nevertheless, the computational cost of performing a forward and a backward pass through a network depends on the its architecture. In this work, we chose to use the same architecture as that used by DDPM [19], in order to isolate the impact of our algorithmic approach from the choice of the model architecture when comparing to DDMs. However, it is an important future direction to explore architectures that are particularly optimized for CDMs and that alleviate the gap in computational complexity. Such architectures should have the property that performing a forward pass and a backward pass through them is computationally similar to performing only a forward pass in a regular DDM. This could potentially be achieved e.g., by relying only the encoder part of the UNet. However, we leave this exploration for future work.
@@ Appendix B Proofs
Since that $p_{{\mathbf{x}}_{T+1}}({\bm{x}})=\mathcal{N}(0,\mathbf{I})$, we have that $\nabla_{\bm{x}}\log(p_{{\mathbf{x}}_{T+1}}({\bm{x}}))=-{\bm{x}}$. Therefore, overall we have
@@
As for the left hand side, since ${\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\mathbf{\varepsilon}}_{t}$ and ${\mathbf{\varepsilon}}_{t}\sim\mathcal{N}(0,\mathbf{I})$, using Tweedie’s formula [33, 44, 11] it can be shown that
@@
B.2 Proof of Tweedie’s Formula for the Variance Preserving Case
Let us show that if ${\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\mathbf{\varepsilon}}_{t}$ and ${\mathbf{\varepsilon}}_{t}\sim\mathcal{N}(0,\mathbf{I})$ then $\nabla_{\bm{x}}\log(p_{{\mathbf{x}}_{t}}({\bm{x}}))=-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\mathbb{E}[{\mathbf{\varepsilon}}_{t}|{\mathbf{x}}_{t}={\bm{x}}]$.
Note that in this case
@@ B.2 Proof of Tweedie’s Formula for the Variance Preserving Case
Taking the gradient w.r.t. ${\bm{x}}_{t}$ gives
$\nabla_{{\bm{x}}_{t}}p_{{\mathbf{x}}_{t}}({\bm{x}}_{t})$ $=\int-\frac{1}{1-\bar{\alpha}_{t}}({\bm{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0})\frac{1}{(2\pi)^{d/2}(1-\bar{\alpha}_{t})^{d/2}}\cdot$
$\quad\qquad\exp\left\{-\frac{1}{2(1-\bar{\alpha}_{t})}\lVert{\bm{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}\rVert^{2}\right\}p_{{\mathbf{x}}_{0}}({\bm{x}}_{0})d{\bm{x}}_{0}$
$=\int-\frac{1}{1-\bar{\alpha}_{t}}({\bm{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0})p_{{\mathbf{x}}_{t}|{\mathbf{x}}_{0}}({\bm{x}}_{t}|{\bm{x}}_{0})p_{{\mathbf{x}}_{0}}({\bm{x}}_{0})d{\bm{x}}_{0}$
$=\int-\frac{1}{1-\bar{\alpha}_{t}}({\bm{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0})p_{{\mathbf{x}}_{0}|{\mathbf{x}}_{t}}({\bm{x}}_{0}|{\bm{x}}_{t})p_{{\mathbf{x}}_{t}}({\bm{x}}_{t})d{\bm{x}}_{0}.$ (20)
Dividing both sides by $p_{{\mathbf{x}}_{t}}({\bm{x}}_{t})$, we get
$\nabla_{{\bm{x}}_{t}}\log p_{{\mathbf{x}}_{t}}({\bm{x}}_{t})$ $=\int-\frac{1}{1-\bar{\alpha}_{t}}({\bm{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0})\,p_{{\mathbf{x}}_{0}|{\mathbf{x}}_{t}}({\bm{x}}_{0}|{\bm{x}}_{t})d{\bm{x}}_{0}$
$=\int-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\,{\mathbf{\varepsilon}}_{t}\,p_{{\mathbf{x}}_{0}|{\mathbf{x}}_{t}}({\bm{x}}_{0}|{\bm{x}}_{t})d{\bm{x}}_{0}$
$=-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\mathbb{E}[{\mathbf{\varepsilon}}_{t}|{\mathbf{x}}_{t}={\bm{x}}_{t}],$ (21)
completing the proof.

### v3

@@ 4 Experiments
We train several CDMs on two common datasets. For CIFAR-10 [26] we train both a class conditional model and an unconditional model. We also train a similar model for CelebA [31], using face images of size $64\times 64$. In Sec. 4.1, we demonstrate why existing DRE methods fail on complex high-dimensional data like images, and show how the incorporation of the MSE loss in our method, according to Theorem 3.1, overcomes these challenges. In Sec. 4.2, we compare our method with pre-trained DDMs with similar architectures, to disentangle the benefits of our method from other variables. We evaluate the performance of CDM as a denoiser, assess its generation quality using FID [17], and measure its likelihood modeling capabilities using NLL. Finally, in Sec. 4.3, we demonstrate the use of different noise schedulers, one specifically tuned for likelihood estimation, and one corresponding to the flow matching optimal-transport scheme [29, 30]. We show that incorporating these schedulers into our method leads to state-of-the-art NLL results among methods capable of outputting the likelihood in a single forward pass.
@@ 4.1 The Importance of Using Both Losses for Achieving an Optimal Classifier
In Theorem 3.1, we established that the MMSE denoiser corresponds to the gradient of the optimal noise-level classifier. A natural question is whether we can train our model only with the MSE loss. Unfortunately, the answer is negative. This is because the MSE achieved by the model does not change if we add a function of $t$ to its output, as such an additive term vanishes when taking the gradient with respect to ${\bm{x}}$. The CE loss is important for removing this degree of freedom. Namely, without the CE loss, the model can function as a denoiser but is useless for the purpose of outputing the likelihood in a single step.
Can we train the model only with the CE loss, then? In theory, training the model only with the CE loss should be sufficient. However, as we will demonstrate next, incorporating MSE is crucial in practice for achieving an optimal classifier.
Table 1 reports the MSE, CE and classification accuracy achieved by models trained with different losses. We emphasize that the model trained using only MSE in this comparison, is a CDM model trained using Algorithm 1, and is not equivalent to a DDM model trained using (5). As evident from the table, when using only the CE loss, the MSE is high, and when using only the MSE loss the CE and classification accuracy are poor. An important point to notice is that even when training with the CE loss, the classifier’s accuracy is rather low (though greater than the $0\%$ achieved when training only with the MSE loss). This is a key prerequisite for making DRE methods work. Specifically, as shown in [35], the classification problem should be sufficiently hard in order to avoid the density-chasm problem, otherwise the classifier can easily discriminate between the classes even without having learned the correct density ratio. Yet, as we illustrate next, only making the classification problem harder is still insufficient for learning the probability $p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})$ with only the CE loss.
Figure 3 shows the logits $f_{\theta}({\bm{x}}_{t})$ for noisy images with different noise levels, comparing a model trained using CE to a model trained with both CE and MSE. As can be seen, in both scenarios the prediction near the true label is the same, namely the CE works well in the vicinity of the correct noise level. However, the model trained without the MSE loss exhibits significantly higher predicted logits for more distant noise levels compared to the model trained using both CE and MSE. Moreover, the logits of the model trained without MSE do not decrease monotonically as the distance from the actual noise level increases, which is in contrast with the expected behavior. This demonstrates the importance of the MSE loss for obtaining good prediction globally. As can be seen in Theorem 3.1, the denoiser at timestep $t$ depends on the predictions of the classifier in both the $t$-th and the $(T+1)$-th entries. Therefore, the addition of the MSE loss enforces the classifier to achieve accurate predictions in both entries, thereby ensuring accurate predictions globally.
@@ 4.2 Denoising Results, Image Quality and Negative Log Likelihood
We compare our method with pre-trained DDMs of similar architectures. Since CDM is a classifier and DDM is a timestep-conditional denoiser, we take the architecture of our CDM to be identical to the DDM, except for altering the last two layers to output a vector of logits, and removing all timestep conditioning layers. These changes have a negligible effect on the number of parameters in the model. For more details please refer to App. C.
@@
As shown in Fig. 5, the denoising performance of our CDM surpasses that of pre-trained DDMs at high noise levels, and is comparable to them at lower noise levels. These quantitative results are corroborated by the qualitative examples in Fig. 5, which showcase image denoising results across various noise levels.
@@
The good denoising performance of CDM translates into high quality image generation. This is illustrated qualitatively in Fig. 1, which shows samples from models trained on CelebA and on CIFAR-10 (unconditional). To quantitatively compare the generation quality of CDM to that of pre-trained DDMs, we use 50k FID [17] against the train-set. For both CDMs and DDMs, we compare images sampled using the DDPM [19], DDIM [39], and DPM-Solver (DPMS) [32] samplers, using 1000, 50, and 25 timesteps, respectively. The results, shown in Table 2, demonstrate that CDM is at least comparable to pre-trained DDMs in image quality, outperforming them in most cases.
@@
Figure 4: Denoising performance. The plots show the MSEs (top) and the ratio between the MSEs (bottom) achieved by a pre-trained DDM and by a CDM with the same architecture, as a function of the noise level (timestep $t$). The CDM significantly outperforms the pre-trained DDM at high noise levels, while demonstrating comparable performance at lower noise levels.
@@
Figure 5: Denoising results. The figure depicts a comparison between denoising results on the CelebA dataset for several different noise levels, obtained with a CDM and with a pre-trained DDM with the same architecture. The right column shows the models’ predictions for pure Gaussian noise, which should theoretically be the expectation of the prior distribution. As observed, DDM outputs a highly noisy image, whereas CDM generates an image much closer to the mean of the dataset.
@@
Additionally, we evaluate CDM’s effectiveness in applying classifier-free guidance (CFG) [18] for conditional sampling tasks. As expected, incorporating CFG improves image quality beyond unconditional generation, as reflected in the conditional CIFAR-10 FID results of Table 2. More details and qualitative results are provided in Appendix C.5. These results showcase the effectiveness of CDM for image generation, showing it to be equal or better than a similar DDM.
@@
Finally, we calculate log-likelihoods and compare our NLL results to recent leading methods in Table 3. CDM demonstrates comparable performance on NLL estimation for CIFAR-10 compared to DDMs. Notably, CDM stands out as a more efficient method than existing ones, requiring only a single forward pass for NLL computation. Table 3 also includes CDM(unif.), and CDM(OT) on which we elaborate in Sec. 4.3 below. These variants of our model, improve the NLL predictions and achieve state-of-the-art results among methods requiring a single step.
4.3 Different Noise Scheduling for Better Likelihood Estimation
@@
To see the effect of using a timestep scheduler tuned for DRE tasks, we repeat the CIFAR-10 unconditional experiment, with a different noise scheduler. Following Rhodes et al. [35] we use the scheduler defined as $\sqrt{1-\bar{\alpha}_{t}}=\frac{t}{T+1}$ and choose $T=1000$ similarly to our previous experiments. Utilizing this scheduler, we achieve a better NLL of 2.98 at the expense of a higher FID of 10.28, when using the DDIM sampler with 50 steps. This trade-off highlights that the scheduler optimal for learning the data distribution may not be ideal for sampling.
To further explore the influence of the noise scheduler, we train and evaluate a CDM with the flow-matching optimal-transport (OT) scheduler [29, 30] on unconditional CIFAR-10. In this scheduler, ${\mathbf{x}}_{t}=\frac{T-t}{T}{\mathbf{x}}_{0}+\frac{t}{T}{\mathbf{\varepsilon}}_{t}$, where ${\mathbf{\varepsilon}}_{t}\sim\mathcal{N}(0,\mathbf{I})$ and $t\in\{0,\ldots,T\}$. This scheduler leads to a state-of-the-art single-step NLL of 2.89 and to an FID of 7.07 with 1000 sampling steps. Please see App. B.4 for more details.
Future research could explore schedulers aimed at further enhancing the NLL. Further analysis of the difference between the schedulers from a classification perspective can be found in App. E.2
@@ 5 Related Work
Using the concept of DRE for learning data distributions was initially studied by Gutmann and Hyvärinen [15]. Their noise contrastive estimation (NCE) method approximates the ratio between the density of the data distribution and that of white Gaussian noise. However, it struggles in practical scenarios where the gap between these distributions is large, as is the case for natural images [35]. Conditional noise contrastive estimation (CNCE) [4] is a slightly improved version of NCE, in which the classification problem is designed to be harder. Specifically, CNCE is based on training a classifier to predict the order of a pair of samples with closer densities, e.g. achieved by pairing a data sample with its noisy version.
Telescoping density-ratio estimation (TRE), proposed by Rhodes et al. [35], avoids direct classification between data and noise. Instead, it uses a gradual transition between those two distributions, and trains a classifier to distinguish between samples from every pair of adjacent densities. Such a classifier learns the ratio between adjacent distributions, and the overall ratio between the data and noise distributions is computed by multiplying all intermediate ratios.
Choi et al. [7] extended this idea from a finite set of intermediate densities to an infinite continuum. This was accomplished by deriving a link between the density ratios for infinitesimally close distributions and the principles of score matching [23, 40, 42], motivating the training of a model to predict the time score $\frac{\partial}{\partial t}\log p_{{\mathbf{x}}_{t}}$. In contrast, we draw a different connection which shows that an MMSE denoiser can be obtained as the gradient of an optimal noise level classifier. Also, to obtain the log ratio between the target and reference distributions, Choi et al. [7] need to solve an integral over the time-score using an ODE solver, while in our method this ratio can be calculated in a single NFE.
@@
Table 2: Image generation quality. We compare the FID (lower is better) achieved by a DDM and a CDM using three sampling schemes for CelebA and CIFAR-10. For conditional CIFAR-10 we train a DDM ourselves, as no model in the original implementation [19] supports CFG.
@@
Table 3: NLL (bits/dim) calculated on the CIFAR-10 test-set. For each model we specify the number of NFEs required for calculating the NLL. CDM achieves state-of-the-art NLL among methods that use a single NFE.
@@ 6 Discussion and Conclusion
We developed an analytical connection between an MSE-optimal denoiser for removing white Gaussian noise and a cross-entropy-optimal classifier for predicting the noise level. We used this connection to propose CDM – a DRE based generative technique that is based on a noise-level classifier. Importantly, our classifier is trained using both a classification loss (CE) and a regression loss (MSE). We showed that this key component is what sets CDM apart from existing DRE based methods, and makes it the first instance of a DRE-based technique that can successfully generate images beyond MNIST.
Our approach is not free of limitations. A key challenge is that CDMs can be more computationally expensive than DDMs. Indeed, while DDMs require a single forward pass for each denoising step, CDMs require both a forward pass and a backward pass. Nevertheless, the computational cost of performing a forward and a backward pass through a network depends on its architecture. In this work, we chose to use the same architecture as that used by DDPM [19], in order to isolate the impact of our algorithmic approach from the choice of the model architecture when comparing to DDMs. However, an important future direction would be to explore architectures that are particularly optimized for CDMs and that alleviate the gap in computational complexity. Such architectures should have the property that performing a forward pass and a backward pass through them is computationally similar to performing only a forward pass in a regular DDM. This could potentially be achieved e.g., by relying only on the encoder part of the U-Net. However, we leave this exploration for future work.
@@
Acknowledgments
This research was partially supported by the Israel Science Foundation (ISF) under Grant 2318/22.
@@ Appendix B Proofs
Since $p_{{\mathbf{x}}_{T+1}}({\bm{x}})=\mathcal{N}({\bm{x}};0,\mathbf{I})$, we have that $\nabla_{\bm{x}}\log(p_{{\mathbf{x}}_{T+1}}({\bm{x}}))=-{\bm{x}}$. Therefore, overall we have
@@
As for the left hand side of (15), since ${\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\mathbf{\varepsilon}}_{t}$ and ${\mathbf{\varepsilon}}_{t}\sim\mathcal{N}(0,\mathbf{I})$, using Tweedie’s formula [34, 45, 11] it can be shown that
@@
B.2 Proof of Tweedie’s Formula
Let us show that if ${\mathbf{y}}=\mu{\mathbf{x}}+\sigma{\mathbf{z}}$ with ${\mathbf{z}}\sim\mathcal{N}(0,\mathbf{I})$ statistically independent of ${\mathbf{x}}$, then $\nabla_{\bm{y}}\log(p_{{\mathbf{y}}}({\bm{y}}))=-\frac{1}{\sigma}\mathbb{E}[{\mathbf{z}}|{\mathbf{y}}={\bm{y}}]$.
From the law of total probability,
@@ B.2 Proof of Tweedie’s Formula
Taking the gradient w.r.t. ${\bm{y}}$ gives
$\nabla_{{\bm{y}}}p_{{\mathbf{y}}}({\bm{y}})$ $=\int-\frac{1}{\sigma^{2}}({\bm{y}}-\mu{\bm{x}})\frac{1}{(2\pi)^{d/2}\sigma^{d}}\exp\left\{-\frac{1}{2\sigma^{2}}\lVert{\bm{y}}-\mu{\bm{x}}\rVert^{2}\right\}p_{{\mathbf{x}}}({\bm{x}})d{\bm{x}}$
$=\int-\frac{1}{\sigma^{2}}({\bm{y}}-\mu{\bm{x}})p_{{\mathbf{y}}|{\mathbf{x}}}({\bm{y}}|{\bm{x}})p_{{\mathbf{x}}}({\bm{x}})d{\bm{x}}$
$=\int-\frac{1}{\sigma^{2}}({\bm{y}}-\mu{\bm{x}})p_{{\mathbf{x}}|{\mathbf{y}}}({\bm{x}}|{\bm{y}})p_{{\mathbf{y}}}({\bm{y}})d{\bm{x}}.$ (20)
Dividing both sides by $p_{{\mathbf{y}}}({\bm{y}})$, we get
$\nabla_{{\bm{y}}}\log p_{{\mathbf{y}}}({\bm{y}})$ $=\int-\frac{1}{\sigma^{2}}({\bm{y}}-\mu{\bm{x}})\,p_{{\mathbf{x}}|{\mathbf{y}}}({\bm{x}}|{\bm{y}})d{\bm{x}}$
$=-\frac{1}{\sigma^{2}}\mathbb{E}[{\mathbf{y}}-\mu{\mathbf{x}}|{\mathbf{y}}={\bm{y}}]$
$=-\frac{1}{\sigma}\mathbb{E}[{\mathbf{z}}|{\mathbf{y}}={\bm{y}}].$ (21)
Now, substituting $\mu=\sqrt{\bar{\alpha}_{t}}$, $\sigma=\sqrt{1-\bar{\alpha}_{t}}$, ${\mathbf{y}}={\mathbf{x}}_{t}$, ${\bm{y}}={\bm{x}}_{t}$, and ${\mathbf{z}}=\varepsilon_{t}$, leads to
$\nabla_{{\bm{x}}_{t}}\log p_{{\mathbf{x}}_{t}}({\bm{x}}_{t})=-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\mathbb{E}[{\mathbf{\varepsilon}}_{t}|{\mathbf{x}}_{t}={\bm{x}}_{t}],$ (22)
demonstrating (17).
@@
B.4 Proof of CDM for the Flow Matching Optimal Transport Scheduler
In the case of the flow matching optimal-transport scheduler [29, 30], ${\mathbf{x}}_{t}$ is defined as
${\mathbf{x}}_{t}=\frac{T-t}{T}{\mathbf{x}}_{0}+\frac{t}{T}{\mathbf{\varepsilon}}_{t},$ (24)
where ${\mathbf{\varepsilon}}_{t}\sim\mathcal{N}(0,\mathbf{I})$ and $t\in\{0,\ldots,T\}$. In this case, the objective of conditional flow matching is [29, 30]
$\mathcal{L}=\sum_{t=1}^{T}\mathbb{E}_{{\mathbf{x}}_{0},\varepsilon_{t}}\left[v_{t}({\bm{x}})-\frac{1}{T}(\varepsilon_{t}-{\mathbf{x}}_{0})\right]$ (25)
and the global minimum of this objective is achieved by
$v_{t}({\bm{x}})=\frac{1}{T}\cdot\mathbb{E}[\varepsilon_{t}-{\mathbf{x}}_{0}|{\mathbf{x}}_{t}={\bm{x}}].$ (26)
We will start by expressing this solution in terms of the optimal denoiser $\mathbb{E}[\varepsilon_{t}|{\mathbf{x}}_{t}]$. Substituting (24) into (26) and using the fact that $\mathbb{E}[{\mathbf{x}}_{t}|{\mathbf{x}}_{t}={\bm{x}}]={\bm{x}}$ gives
$v_{t}({\bm{x}})=\frac{1}{T-t}\cdot(\mathbb{E}[\varepsilon_{t}|{\mathbf{x}}_{t}={\bm{x}}]-{\bm{x}}).$ (27)
Next, following exactly the same logic as in App. B.1, we can write
$\mathbb{E}[\varepsilon_{t}|{\mathbf{x}}_{t}={\bm{x}}]=\frac{t}{T}\cdot\left(\nabla_{{\bm{x}}}\log(p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(T|{\bm{x}}))-\nabla_{{\bm{x}}}\log(p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}}))+{\bm{x}}\right),$ (28)
Finally, by substituting (28) into (27) we get
$v_{t}({\bm{x}})=\frac{t/T}{T(1-t/T)}\left(\nabla_{{\bm{x}}}\log(p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(T|{\bm{x}}))-\nabla_{{\bm{x}}}\log(p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}}))\right)-\frac{1}{T}\cdot{\bm{x}}.$ (29)